# VoxEval: Benchmarking the Knowledge Understanding Capabilities of End-to-End Spoken Language Models

### v2

@@ Abstract
With the growing demand for developing speech-based interaction models, end-to-end Spoken Language Models (SLMs) have emerged as a promising solution. When engaging in conversations with humans, it is essential for these models to comprehend a wide range of world knowledge. In this paper, we introduce VoxEval, a novel speech question-answering benchmark specifically designed to assess SLMs’ knowledge understanding through purely speech-based interactions. Unlike existing AudioQA benchmarks, VoxEval maintains speech format for both questions and answers, evaluates model robustness across diverse audio conditions (varying timbres, audio qualities, and speaking styles), and pioneers the assessment of challenging domains like mathematical problem-solving in spoken format. Our comprehensive evaluation of recent SLMs using VoxEval reveals significant performance limitations in current models, highlighting crucial areas for future improvements. [^1] VoxEval dataset is available at: https://github.com/dreamtheater123/VoxEval
@@
Wenqian Cui1, Xiaoqi Jiao2, Ziqiao Meng3, Irwin King1, 1The Chinese University of Hong Kong, 2LightSpeed Studios, Tencent, 3National University of Singapore
@@ 1 Introduction
The progress in text-based Large Language Models (LLMs) has significantly impacted the field of generative AI, allowing for smooth communication between humans and AI via written text. However, since natural human interactions often involve spoken language, there is a growing movement to integrate speech capabilities into LLMs Cui et al. (2024). This initiative aims to enable direct speech communication between humans and AI by endowing language models with the ability to understand and produce speech. Consequently, end-to-end Spoken Language Models (SLMs), which engage directly in spoken dialogue, represent an exciting and promising direction.
Figure 1: Illustration of the three crucial factors for evaluating SLMs’ knowledge comprehension.
Evaluation of SLMs is crucial for ensuring their effectiveness and reliability in real-world applications. Similar to Text Language Models (TLM), the ability to understand world knowledge is essential for SLMs to perform conversations with humans. When evaluating SLMs’ knowledge comprehension, several crucial factors should be taken into account: 1) The evaluation should be conducted entirely in speech format, meaning both the input and output must be audio-based. This approach differs from typical AudioQA tasks Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024), where audio clips are matched with text questions. SLM evaluation should solely use speech to mimic real-world conversational settings. 2) The evaluation should consider various input audio conditions. Users may interact with SLMs under diverse audio characteristics, such as different speaker timbres, audio qualities, and speaking styles. Consequently, SLMs should demonstrate robustness against these varying conditions, producing consistent answers when the input content remains the same OpenAI (2024). 3) Not all knowledge comprehension topics are straightforward for SLM evaluation. The range of knowledge understanding includes diverse subjects, including literature, sociology, engineering, etc. However, some subjects present unique challenges when evaluated in the spoken format. For instance, while math expressions can be easily and concisely written in text, evaluating them in spoken form remains a challenge. Figure 1 illustrates the three key elements mentioned above.
Based on the key elements outlined, we introduce VoxEval, a speech question-answering benchmark designed to assess the knowledge understanding capabilities of SLMs. This benchmark features audio recordings for both questions and instructions, facilitating a comprehensive evaluation of speech modeling abilities. Additionally, VoxEval incorporates a variety of input conditions for the same question, allowing us to test the robustness of SLMs across different audio scenarios. To address the challenges associated with evaluating certain subjects, such as mathematics, we pinpoint the critical differences between written and spoken formats and convert the former into their spoken counterparts. Our evaluation of recently proposed SLMs using VoxEval reveals that it presents a significant challenge for existing models. Additionally, existing SLMs demonstrate vulnerability to changes in input audio conditions, indicating the need for future research to focus on improving robustness. To summarize our contributions:
@@
  • We introduce a novel benchmark, VoxEval, specifically designed to evaluate the knowledge understanding capabilities of SLMs exclusively in speech format, across a diverse range of audio conditions.
@@
  • We pioneer the assessment of SLMs’ math-solving abilities, addressing a critical gap in existing evaluation methodologies.
@@
  • We conduct a systematic evaluation of VoxEval with several recent SLMs, revealing that current models exhibit significant limitations in performance on our benchmark.
@@ 2.2 SpeechLLM Evaluation Benchmarks
The evaluation of SpeechLLMs centers on their ability to model speech data, with researchers examining various aspects and levels. Initially, multiple datasets are utilized to assess the linguistic modeling capabilities of SpeechLLMs. For instance, sWUGGY Nguyen et al. (2020) is employed for lexical evaluation, sBLIMP Nguyen et al. (2020) for syntactical assessment, and sStoryCloze Hassid et al. (2024) for semantic evaluation. Additionally, the modeling of paralinguistic elements (spoken content beyond words) is crucial for SpeechLLMs, prompting the development of datasets that evaluate paralinguistic information from both token and perceptual levels Kharitonov et al. (2022); Nguyen et al. (2024). Beyond the modeling aspect, the performance of downstream applications is vital for the practical use of SpeechLLMs in real-world scenarios. Recently, numerous downstream evaluation benchmarks have emerged Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024), focusing on general audio understanding, including speech, environmental sounds, and music. However, as highlighted in Section 1, these benchmarks contain all QA pairs in text format, making them suitable for evaluating S2TLLMs but not SLMs. The closest benchmark to our work is VoiceBench Chen et al. (2024), which also evaluates end-to-end SLMs under various audio conditions. However, our study places a stronger emphasis on evaluating the knowledge comprehension abilities of SLMs, not only by including a broader range of topics but also by particular focusing on their complex skills like mathematics. We also offer a detailed analysis of the performance of SLMs across various scenarios.
@@ 3 VoxEval
Filler Words Include words like "um," "uh," "like," and "you know" to see if the model can accurately process the main content despite interruptions. Uh, I’m not sure if I can, like, make it to the meeting on time.
@@ 3.1 Data Construction
  • MMLU is widely used for evaluating SLMs Nguyen et al. (2024); Défossez et al. (2024). However, this evaluation primarily reflects their capability to handle text, not speech.
@@
To transform MMLU into its spoken version, we need to concatenate all the text within a question into a single sequence. In MMLU, every data is a Multiple Choice Question (MCQ) with four answer choices. We use a simple sentence to concatenate the question and answer choices. Therefore, every converted data looks like this:
@@
Among the 57 subjects in MMLU, we exclude the "high school computer science" subject when constructing VoxEval, as it includes code snippets that are not suitable for verbal evaluation.
@@ 3.2.2 Different Speaking Styles
Linguistic variation involves changes in the content. For instance, people may sometimes pause to think about what they want to say, leading to disfluencies in their speech. Another example is when someone accidentally says something incorrectly and then corrects themselves upon realizing the mistake. SLMs should still accurately interpret the speech in these situations. However, these scenarios are not typically captured when using TTS to generate speech from a well-written text. To address this, we propose five techniques for linguistic variation to accommodate different situations, including filler words, mispronunciations, disfluencies, false starts and corrections, and language proficiency. Table 1 provides explanations and examples for each of these five variations. To replicate these scenarios, we first use GPT-4o OpenAI (2024) to modify the original question text into different versions and then convert them into speech. We only alter the question text and not the answer choice texts.
@@
Alloy 0.0001 0.0480 0.2084 0.1216 0.3763
Echo 0.0001 0.0558 0.2096 0.1221 0.3764
Fable 0.0000 0.0116 0.2084 0.1153 0.3642
Nova 0.0001 0.0332 0.2070 0.1298 0.3677
Onyx 0.0002 0.0275 0.1966 0.1192 0.3764
Shimmer 0.0000 0.0516 0.2076 0.1205 0.3815
@@
Linguistic 0.0001 0.0488 0.2044 0.1187 0.3643
Speed 0.0001 0.0503 0.1911 0.1013 0.3469
Pitch 0.0000 0.0544 0.1788 0.0609 0.3345
@@
Noise 0.0000 0.0368 0.1950 0.1018 0.3695
Other Env Acoustics 0.0001 0.0434 0.2019 0.1051 0.3728
Underlying Text LMs Llama-7B Llama-7B Llama-2-7B Helium-7B GLM-4-9B
Text MMLU 0.3510 0.3510 0.4530 0.5430 0.7470
Table 3: VoxEval evaluation results of popular SLMs under various input audio conditions.
Paralinguistic variation, in contrast, involves changes other than the contents of the input audio, without changing the content. Similar to linguistic variation, SLMs should also output consistent outputs under different paralinguistic variations. We consider four types of paralinguistic variations.
  • Pitch shift: Pitch shift refers to the alteration of the pitch of the speech signal, which can simulate the effect of someone speaking in a higher or lower tone. For each question input, we randomly shift its pitch between -5 and 5 semitones.
  • Tempo change: In real-life situations, people often vary their speaking speed, sometimes talking faster and other times slower. To mimic these variations in speaking pace, we apply tempo changes. For each question input, we randomly select a tempo-changing rate that ranges from half the original speed to twice as fast.
@@ 3.2.3 Different Audio Qualities
  • Background noise: Sometimes the input audio may be accompanied by background noise, which can affect the clarity and intelligibility of the spoken content. We consider different kinds of noise, including Gaussian noise, colored noise, background music, and short interruption noise. For each question input, a kind of noise is randomly selected and applied to the original audio.
  • Other environment acoustics: We consider speech recorded under various conditions, including aliasing, room impulse response, low pass filter, high pass filter, band pass filter, bitcrush, gain, clipping distortion, and seven band parameter EQ. Full explanations and transformation settings are presented in Appendix A. A random audio effect is applied to each question input.
@@
We apply all the variations to the speech data generated only using the speaker "alloy". To create different paralinguistic and acoustic conditions, we utilize the “audioaugmentations" package [^3] https://github.com/iver56/audiomentations to augment the audio.
@@ 3.3 Handle Math Expressions and More
SLMs (MGS) SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice
Speakers
Alloy 0.0000/0.0000/0.0000/0.0003 0.0561/0.0381/0.0319/0.0587 0.2244/0.2171/0.1929/0.2330 0.1309/0.1195/0.1198/0.1285 0.3250/0.3858/0.3973/0.4303
Echo 0.0004/0.0000/0.0000/0.0000 0.0611/0.0548/0.0341/0.0692 0.2248/0.2182/0.1867/0.2422 0.1413/0.1246/0.1129/0.1371 0.3325/0.4226/0.3906/0.4123
Fable 0.0000/0.0000/0.0000/0.0000 0.0088/0.0056/0.0044/0.0170 0.2122/0.2290/0.1915/0.2364 0.1257/0.1165/0.1042/0.1191 0.3081/0.4252/0.3832/0.3933
Nova 0.0000/0.0002/0.0000/0.0001 0.0349/0.0274/0.0197/0.0408 0.2302/0.2172/0.1799/0.2291 0.1504/0.1269/0.1084/0.1432 0.3195/0.4280/0.3843/0.4012
Onyx 0.0001/0.0000/0.0001/0.0000 0.0346/0.0273/0.0149/0.0395 0.2085/0.2109/0.1823/0.2242 0.1356/0.1274/0.1105/0.1339 0.3238/0.4427/0.4084/0.3832
Shimmer 0.0000/0.0000/0.0000/0.0000 0.0686/0.0479/0.0301/0.0691 0.2190/0.2229/0.1909/0.2172 0.1387/0.1210/0.1085/0.1306 0.3179/0.4596/0.3935/0.4434
Speaking Styles
Linguistic 0.0000/0.0000/0.0004/0.0000 0.0581/0.0431/0.0305/0.0588 0.2062/0.2157/0.2020/0.2257 0.1310/0.1167/0.1009/0.1344 0.3218/0.3818/0.3694/0.3991
Speed 0.0000/0.0000/0.0002/0.0005 0.0571/0.0514/0.0342/0.0667 0.2063/0.2094/0.1783/0.2038 0.1059/0.1137/0.0916/0.1171 0.3055/0.3874/0.3630/0.3798
Pitch 0.0000/0.0000/0.0000/0.0000 0.0620/0.0555/0.0367/0.0685 0.1794/0.1844/0.1760/0.2019 0.0660/0.0573/0.0567/0.0585 0.3003/0.3735/0.3476/0.3643
Audio Qualities
Noise 0.0000/0.0000/0.0000/0.0000 0.0422/0.0337/0.0247/0.0484 0.2077/0.2173/0.1659/0.2149 0.1073/0.1079/0.0983/0.1078 0.3145/0.4193/0.3868/0.4146
Other Env Acoustics 0.0000/0.0000/0.0000/0.0003 0.0519/0.0371/0.0279/0.0546 0.2244/0.2066/0.1684/0.2325 0.1118/0.1008/0.1024/0.1130 0.3045/0.4340/0.3909/0.4142
Underlying Text LMs Llama-7B Llama-7B Llama-2-7B Helium-7B GLM-4-9B
Text MMLU 0.3050/0.3830/0.3400/0.3810 0.3050/0.3830/0.3400/0.3810 - - -
Table 4: Average scores for various subject groups across different input audio conditions are presented. The scores are shown in the format “STEM/Social Science/Humanities/Others" subjects. MGD stands for Mean Grouped Score. The highest score in each scenario is highlighted in bold.
@@
SLMs Elementary Math High School Math College Math
SpeechGPT 0.0000 0.0000 0.0000
TWIST 0.0317 0.0444 0.0800
SPIRIT-LM 0.2302 0.2000 0.2200
Moshi 0.1138 0.1926 0.1700
GLM-4-Voice 0.2566 0.2593 0.2700
Table 5: Evaluation results for different levels of math problems under the “Alloy" speaker setting.

### v3

@@ Abstract
With the rising need for speech-based interaction models, end-to-end Spoken Language Models (SLMs) have emerged as a promising solution. While these models require comprehensive world knowledge for meaningful and reliable human interactions, existing question-answering (QA) benchmarks fall short in evaluating SLMs’ knowledge understanding due to their inability to support end-to-end speech evaluation and account for varied input audio conditions. To address these limitations, we present VoxEval, a novel SpeechQA benchmark that assesses SLMs’ knowledge understanding through pure speech interactions. Our benchmark 1) uniquely maintains speech format for both inputs and outputs, 2) evaluates model robustness across diverse input audio conditions, and 3) pioneers the assessment of complex tasks like mathematical reasoning in spoken format. Systematic evaluation demonstrates that VoxEval presents significant challenges to current SLMs, revealing their sensitivity to varying audio conditions and highlighting the need to enhance reasoning capabilities in future development. We hope this benchmark could guide the advancement of more sophisticated and reliable SLMs. [^1] VoxEval dataset is available at: https://github.com/dreamtheater123/VoxEval
@@
Wenqian Cui1, Xiaoqi Jiao2, Ziqiao Meng3, Irwin King1, 1The Chinese University of Hong Kong, 2LIGHTSPEED STUDIOS, 3National University of Singapore
@@ 1 Introduction
The rapid progress in Text-based Large Language Models (TLMs) has revolutionized generative AI, enabling seamless communication between humans and AI via written text. However, natural human interactions often occur through spoken language, making it essential to extend these capabilities to speech. End-to-end Spoken Language Models (SLMs) have emerged as a promising solution, enabling direct speech-based communication with AI Cui et al. (2024). For SLMs to facilitate meaningful and reliable interactions, it is crucial to evaluate their understanding of world knowledge—an essential capability for effective human-AI dialogue. Such evaluations ensure that SLMs can handle complex conversational scenarios and provide accurate, contextually appropriate responses.
Figure 1: Illustration of the limitations and challenges of SLM knowledge comprehension evaluation.
Existing benchmarks Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024) for evaluating audio-based knowledge comprehension face two major limitations. First, they fail to support end-to-end speech evaluations. These benchmarks typically build datasets by matching audio clips with text-based question/answer pairs, which differ from the speech interaction pipeline. Second, they do not account for diverse input audio conditions, such as variations in speaker timbre, speaking styles, and audio quality, which are common in real-world settings. However, it is crucial for SLMs to demonstrate robustness against these varying conditions, producing consistent answers when the input content remains the same Hurst et al. (2024). Additionally, we identify a unique challenge in SLM knowledge evaluation: not all knowledge comprehension topics are straightforward for SLM evaluation. For example, mathematical reasoning is easy to evaluate in written form due to the concise nature of mathematical expressions. However, evaluating it in spoken form remains a challenge. Despite this, possessing math reasoning skills is crucial for SLMs, as they are essential for applications like tutoring and simple problem-solving. Figure 1 illustrates the three key elements mentioned above.
Based on the key elements outlined, we introduce VoxEval, a SpeechQA benchmark designed to assess the knowledge understanding capabilities of SLMs with the following properties. First, VoxEval includes audio recordings for both questions and answers, enabling end-to-end speech interaction evaluation. Second, it incorporates diverse input audio conditions for the same question, allowing us to test the robustness of SLMs across various audio scenarios. Third, to tackle the challenges of evaluating subjects like mathematics and other complex topics, we identify the critical distinctions between written and spoken formats, transform written content into its spoken equivalent, and incorporate Chain-of-Thought prompts to assess the reasoning ability of SLMs in a step-by-step problem-solving setting. Table 1 presents the statistics for VoxEval, while Table 2 compares the distinctive features of VoxEval with those of other benchmarks. Our evaluation of recently developed SLMs shows that VoxEval poses a substantial challenge for current models, as they perform poorly on VoxEval, exhibit sensitivity to changes in input audio conditions, and demonstrate very limited reasoning capabilities. To summarize our contributions:
@@
  • We introduce a novel benchmark, VoxEval, designed to evaluate the knowledge understanding capabilities of SLMs exclusively in speech format, across a diverse range of audio conditions.
@@
  • We pioneer the assessment of SLMs’ math-solving abilities by introducing a set of spoken math problems and answers.
@@
  • We conduct a systematic evaluation of VoxEval with several recent SLMs, showing VoxEval is a challenging benchmark to existing SLMs.
@@ 2.1 Speech Large Language Models
QA pairs subjects # IAC metric
13,938 56 26 ASR + accuracy
Table 1: Statistics of VoxEval. IAC represents Input Audio Condition.
SLMs e2e # KUPs # variations math CoT
SD-Eval
Dynamic-SUPERB
AIR-Bench
AudioBench 5,196
MMAU $\sim$3,300
VoiceBench 3,074 64,695
VoxEval 13,938 153,318
Table 2: A comparison of our benchmark with other benchmarks. NUP and CoT stand for Knowledge Understanding Problem and Chain-of-Thought prompts.
@@ 2.2 SpeechLLM Evaluation Benchmarks
The evaluation of SpeechLLMs centers on their ability to model speech data, with researchers examining various aspects and levels. Initially, multiple datasets are utilized to assess the linguistic modeling capabilities of SpeechLLMs. For instance, sWUGGY Nguyen et al. (2020) is employed for lexical evaluation, sBLIMP Nguyen et al. (2020) for syntactical assessment, and sStoryCloze Hassid et al. (2024) for semantic evaluation. Additionally, the modeling of paralinguistic elements (spoken content beyond words) is crucial for SpeechLLMs, prompting the development of datasets that evaluate paralinguistic information from both token and perceptual levels Kharitonov et al. (2022); Nguyen et al. (2024). Beyond the modeling aspect, the performance of downstream applications is vital for the practical use of SpeechLLMs in real-world scenarios. Recently, numerous downstream evaluation benchmarks have emerged Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024), focusing on general audio understanding, including speech, environmental sounds, and music. However, as highlighted in Section 1, these benchmarks contain all QA pairs in text format, making them suitable for evaluating S2TLLMs but not SLMs. The closest benchmark to our work is VoiceBench Chen et al. (2024), which also evaluates end-to-end SLMs under various audio conditions. However, our study places a stronger emphasis on evaluating the knowledge comprehension abilities of SLMs, not only by including a broader range of topics but also by particularly focusing on their complex skills like mathematics. We also offer a detailed analysis of the performance of SLMs across various scenarios.
@@ 3 VoxEval
Filler Words Include words like “um," “uh," “like," and “you know" to see if the model can accurately process the main content despite interruptions. Uh, I’m not sure if I can, like, make it to the meeting on time.
@@ 3.1 Data Construction
  • MMLU is widely used for evaluating the additional text processing abilities of SLMs Nguyen et al. (2024); Défossez et al. (2024). However, this evaluation does not accurately reflect the speech-processing abilities of SLMs.
@@
To transform MMLU into its spoken version, we need to concatenate all the text within a problem into a single sequence. In MMLU, every data is a Multiple Choice Question (MCQ) with four answer choices. We use a simple sentence to concatenate the question and answer choices. Therefore, every converted data looks like this:
@@
Among the 57 MMLU subjects, “high school computer science" is excluded from VoxEval due to the code snippets within it being unsuitable for verbal evaluation. As a result, VoxEval contains 13,938 unique SpeechQA pairs.
@@ 3.2.2 Different Speaking Styles
Linguistic variation involves changes in the content. For instance, people may sometimes pause to think about what they want to say, leading to disfluencies in their speech. Another example is when someone accidentally says something incorrectly and then corrects themselves upon realizing the mistake. SLMs should still accurately interpret the speech in these situations. However, these scenarios are not typically captured when using TTS to generate speech from a well-written text. To address this, we propose five techniques for linguistic variation to accommodate different situations, including filler words, mispronunciations, disfluencies, false starts and corrections, and language proficiency. Table 3 provides explanations and examples for each of these five variations. To replicate these scenarios, we first use GPT-4o Hurst et al. (2024) to modify the original question text into different versions and then convert them into speech. We randomly choose a variation technique for each problem. We only alter the question text and not the answer choice texts.
@@
Alloy 0.01 4.80 20.84 12.16 37.63
Echo 0.01 5.58 20.96 12.21 37.64
Fable 0.00 1.16 20.84 11.53 36.42
Nova 0.01 3.32 20.70 12.98 36.77
Onyx 0.02 2.75 19.66 11.92 37.64
Shimmer 0.00 5.16 20.76 12.05 38.15
@@
Linguistic 0.01 4.88 20.44 11.87 36.43
Speed 0.01 5.03 19.11 10.13 34.69
Pitch 0.00 5.44 17.88 6.09 33.45
@@
Noise 0.00 3.68 19.50 10.18 36.95
Other Env Acou 0.01 4.34 20.19 10.51 37.28
Table 5: Evaluation accuracy (%) of popular SLMs in VoxEval evaluations across different input audio conditions.
SLMs $S\rightarrow S$ $S\rightarrow T$ $T\rightarrow S$ $T\rightarrow T$
SPIRIT-LM 20.84 23.77 1.38 17.77
Table 6: Evaluation accuracy (%) of SPIRIT-LM under different modality combinations. $S$ and $T$ represent speech and text, respectively.
Paralinguistic variation, in contrast, involves changes other than the contents of the input audio. Similar to linguistic variation, SLMs should also output consistent outputs under different paralinguistic variations. We consider two types of paralinguistic variations.
  • Pitch shift. Pitch shift refers to the alteration of the pitch of the speech signal, which can simulate the effect of someone speaking in a higher or lower tone. For each question input, we randomly shift its pitch between -5 and 5 semitones.
  • Tempo change. In real-life situations, people often vary their speaking speed, sometimes talking faster and other times slower. To mimic these variations in speaking pace, we apply tempo changes. For each question input, we randomly select a tempo-changing rate that ranges from half the original speed to twice as fast.
@@ 3.2.3 Different Audio Qualities
  • Background noise. Sometimes the input audio may be accompanied by background noise, which can affect the clarity and intelligibility of the spoken content. We consider different kinds of noise, including Gaussian noise, colored noise, background music, and short interruption noise. For each problem, a kind of noise is randomly selected and applied to the original audio.
  • Other environment acoustics. We consider speech recorded under various conditions, including aliasing, room impulse response, low pass filter, high pass filter, band pass filter, bitcrush, gain, clipping distortion, and seven band parameter EQ. Full explanations and transformation settings are presented in Appendix A. A random audio effect is applied to each question input.
@@
We apply all the variations to the speech data generated only using the speaker “alloy". To create different paralinguistic and acoustic conditions, we utilize the “audioaugmentations" package [^3] https://github.com/iver56/audiomentations to augment the audio.
SLMs (MGS) SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice
Speakers
Alloy 0.00/0.00/0.00/0.03 5.61/3.81/3.19/5.87 22.44/21.71/19.29/23.30 13.09/11.95/11.98/12.85 32.50/38.58/39.73/43.03
Echo 0.04/0.00/0.00/0.00 6.11/5.48/3.41/6.92 22.48/21.82/18.67/24.22 14.13/12.46/11.29/13.71 33.25/42.26/39.06/41.23
Fable 0.00/0.00/0.00/0.00 0.88/0.56/0.44/1.70 21.22/22.90/19.15/23.64 12.57/11.65/10.42/11.91 30.81/42.52/38.32/39.33
Nova 0.00/0.02/0.00/0.01 3.49/2.74/1.97/4.08 23.02/21.72/17.99/22.91 15.04/12.69/10.84/14.32 31.95/42.80/38.43/40.12
Onyx 0.01/0.00/0.01/0.00 3.46/2.73/1.49/3.95 20.85/21.09/18.23/22.42 13.56/12.74/11.05/13.39 32.38/44.27/40.84/38.32
Shimmer 0.00/0.00/0.00/0.00 6.86/4.79/3.01/6.91 21.90/22.29/19.09/21.72 13.87/12.10/10.85/13.06 31.79/45.96/39.35/44.34
Speaking Styles
Linguistic 0.00/0.00/0.04/0.00 5.81/4.31/3.05/5.88 20.62/21.57/20.20/22.57 13.10/11.67/10.09/13.44 32.18/38.18/36.94/39.91
Speed 0.00/0.00/0.02/0.05 5.71/5.14/3.42/6.67 20.63/20.94/17.83/20.38 10.59/11.37/09.16/11.71 30.55/38.74/36.30/37.98
Pitch 0.00/0.00/0.00/0.00 6.20/5.55/3.67/6.85 17.94/18.44/17.60/20.19 06.60/05.73/05.67/05.85 30.03/37.35/34.76/36.43
Audio Qualities
Noise 0.00/0.00/0.00/0.00 4.22/3.37/2.47/4.84 20.77/21.73/16.59/21.49 10.73/10.79/9.83/10.78 31.45/41.93/38.68/41.46
Other Env Acoustics 0.00/0.00/0.00/0.03 5.19/3.71/2.79/5.46 22.44/20.66/16.84/23.25 11.18/10.08/10.24/11.30 30.45/43.40/39.09/41.42
Table 7: Average accuracy (%) for various subject groups across different input audio conditions are presented. The scores are shown in the format “STEM/Social Science/Humanities/Others" subjects. MGD stands for Mean Grouped Score. The highest score in each scenario is highlighted in bold.
@@ 3.3 Handle Math Expressions and More
Elementary Math High School Math College Math
SLMs Direct LLM Direct LLM Direct LLM
Direct Answer
SpeechGPT 0.00 0.00 0.00 0.00 0.00 0.00
TWIST 3.17 1.59 4.44 1.11 8.00 4.00
SPIRIT-LM 23.02 20.63 20.00 20.00 22.00 25.00
Moshi 11.38 11.90 19.26 18.52 17.00 14.00
GLM-4-Voice 25.66 25.66 25.93 25.93 27.00 26.00
Chain-of-Thought (Cut)
SpeechGPT 0.00 0.00 0.00 0.00 0.00 0.00
TWIST 2.65 1.32 4.81 3.33 4.00 2.00
SPIRIT-LM 14.29 14.02 16.30 14.44 18.00 14.00
Moshi 12.17 10.32 14.44 11.48 8.00 10.00
GLM-4-Voice 27.25 28.04 14.44 15.93 22.00 22.00
Chain-of-Thought (No cut)
SpeechGPT 0.00 0.00 0.00 0.00 0.00 0.00
TWIST 0.00 0.00 0.00 0.00 0.00 0.00
SPIRIT-LM 15.08 14.02 16.30 15.19 16.00 17.00
Moshi 11.11 11.11 13.33 11.48 15.00 14.00
GLM-4-Voice 27.25 27.78 14.44 15.19 22.00 22.00
Table 8: Evaluation accuracy (%) for different levels of math problems under the “Alloy" speaker setting.